A rectangle has a perimeter of 196 inches. the length is 28 inches more than the width. What is the dimensions?
Answer by jim_thompson5910(35256)

P = 2L + 2W

196 = 2(W+28) + 2W

196 = 2W+56+2W

196 = 56+4W

0 = 56+4W-196

-4W = 56-196

-4W = -140

W = (-140)/(-4)

W = 35

---------------

L = W+28

L = 35+28

L = 63


So the length is 63 inches and the width is 35 inches.